5y = -7 - 2x
2x - 3y = 33

Answer by checkley77(12844)

5y=-7-2x
2x+5y=-7 now multiply this equation by (-1) & add.
2x-3y=33
-2x-5y=7
---------------
-8y=40
y=40/-8
y=-5 answer.
5*-5=-7-2x
-25=-7-2x
2x=-7+25
2x=18
x=18/2
x=9 answer.
proof
2*9-3*-5=33
18+15=33
33=33